According to a recent survey, 50% of people have brown hair, 25% of people have blonde hair, 15% of people have black hair, and 10% of people have red hair. What is the probability of randomly selecting one person who either has blonde hair or black hair?
Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3838)


Answer: 40%

Reason:
Let's say there are 100 people surveyed.
50 have brown hair (since 50% of 100 is 50).
25 have blonde hair
15 have black hair
10 have red hair
Then 25+15 = 40 people have either blonde or black hair out of the 100 total. That's a 40/100 = 40% chance.
The percentage 40% converts to the exact decimal form 0.40 or simply 0.4
If you want to convert to a fraction, then 0.4 = 2/5

Answer by ikleyn(53937)

.

In this problem, you have 4 disjoint subsets of people, that cover the entire set.

THEREFORE, to answer the question, add the percentage for relevant subsets of people.